To solve the ILLP given below by the graphical method :

Maximum Z = 95x₁ + 100x₂

Subject to the constraints

5x₁ + 2x₂ <= 20

x₁ >= 3

x₂ <= 5

x₁ , x₂ are non-negative integers.


1
Expert's answer
2021-05-11T04:10:44-0400

"Z=95X_1+100X_2" (Maximize)


The condition given:

"5X_1+2X_2\leq 20"

"X_1\geq3"

"X_2\leq5"


In the equation "5X_1+2X_2=20"


when we put "X_1=0",then "X_2=10"

When we put "X_2=0",then "X_1=4"


Now when we draw a graph of the line and the given points and shade accordingly then in the feasible region we got three points

that is

A(3,0)

B(4,0)

C(3,2.5)


Now value of Z at A=

"Z_A=95\times3+0=285"


Value of Z at B=

"Z_B=95\times4+0=380"


Value of Z at C=

"Z_C=95\times3+100\times2.5=535"

So the maximum value is achieved at the point C and that value is 535


So we can say that the Maximum of Z is 535 and it occurred at the point (3,2.5)